# A splitting method for numerical relativistic magnetohydrodynamics

### v1

@@ Appendix A Variables conversion
${\cal E}_{(1)}={\cal E}-{\cal E}_{0}=(\mn@boldsymbol{E_{0}}\!\cdot\!\mn@boldsymbol{E_{1}})+(\mn@boldsymbol{B_{(0)}}\!\cdot\!\mn@boldsymbol{B_{(1)}})+\frac{E_{1}^{2}+B_{(1)}^{2}}{2}+w\gamma^{2}-p\,,$ (131)
@@
${S}_{1}=\mn@boldsymbol{S}-\mn@boldsymbol{S}_{0}=\mn@boldsymbol{E_{0}}\!\times\!\mn@boldsymbol{B_{(1)}}+\mn@boldsymbol{E_{1}}\!\times\!\mn@boldsymbol{B_{(0)}}+\mn@boldsymbol{E_{1}}\!\times\!\mn@boldsymbol{B_{(1)}}+w\gamma^{2}\mn@boldsymbol{v}\,,$ (134)
@@
${S}_{0}=\mn@boldsymbol{E_{0}}\!\times\!\mn@boldsymbol{B_{(0)}}\,.$ (136)
@@
${E}_{1}=-\mn@boldsymbol{E}_{0}-\mn@boldsymbol{v}\!\times\!\mn@boldsymbol{B}\,.$ (138)
@@
Next, we use the perfect conductivity condition (138) to eliminate ${E}_{1}$ from the expression (131) for ${\cal E}_{(1)}$. To this end, we first find that
(${E_{0}}\!\cdot\!\mn@boldsymbol{E_{1}})=-E_{0}^{2}-\mn@boldsymbol{E_{0}}\!\cdot\!(\mn@boldsymbol{v}\!\times\!\mn@boldsymbol{B})\,,$
@@
$E_{1}^{2}=E_{0}^{2}+2\mn@boldsymbol{E_{0}}\!\cdot\!(\mn@boldsymbol{v}\!\times\!\mn@boldsymbol{B})+||\mn@boldsymbol{v}\!\times\!\mn@boldsymbol{B}||^{2}\,,$
@@
(${E_{0}}\!\cdot\!\mn@boldsymbol{E_{1}})+\frac{E_{1}^{2}}{2}=-\frac{E_{0}^{2}}{2}+||\mn@boldsymbol{v}\!\times\!\mn@boldsymbol{B}||^{2}\,.$
@@
${\cal E}_{(1)}=\frac{1}{2}(B^{2}v^{2}-(\mn@boldsymbol{v}\!\cdot\!\mn@boldsymbol{B})^{2})+W-p-\frac{E_{0}^{2}}{2}+\frac{B_{(1)}^{2}}{2}+(\mn@boldsymbol{B_{(0)}}\!\cdot\!\mn@boldsymbol{B_{(1)}})\,.$ (144)
@@
¯$\bar{{\cal E}}_{1}={\cal E}_{(1)}+\frac{E_{0}^{2}}{2}-\frac{B_{(1)}^{2}}{2}-(\mn@boldsymbol{B_{(0)}}\!\cdot\!\mn@boldsymbol{B_{(1)}})$ (145)
@@
$a_{2}=\frac{1}{2}(B^{2}v^{2}-E_{0}^{2})-{\cal E}_{(1)}+\frac{B_{(1)}^{2}}{2}+(\mn@boldsymbol{B_{(0)}}\!\cdot\!\mn@boldsymbol{B_{(1)}})+D\frac{(1-v^{2})^{1/2}}{\kappa}\,.$
The first two terms of this expression constitute the difference between $B^{2}v^{2}/2$ and $E_{0}^{2}$. These non-negative terms can be very large and their difference can be a source of large error in computations of $a_{2}$ in the case of high magnetisation.
@@
${v}_{0}=\frac{\mn@boldsymbol{E_{0}}\!\times\!\mn@boldsymbol{B_{(0)}}}{B_{(0)}^{2}}\,.$
@@
$B^{2}v^{2}-E_{0}^{2}=B^{2}v^{2}-B_{(0)}^{2}v_{0}^{2}=B_{(0)}^{2}(v^{2}-v_{0}^{2})+(B_{(1)}^{2}+(\mn@boldsymbol{B_{(0)}}\!\cdot\!\mn@boldsymbol{B_{(1)}}))v^{2}\,,$
@@
$a_{2}=\frac{1}{2}(B_{(0)}^{2}(v^{2}-v_{0}^{2})+(B_{(1)}^{2}+(\mn@boldsymbol{B_{(0)}}\!\cdot\!\mn@boldsymbol{B_{(1)}}))v^{2})-{\cal E}_{(1)}+\frac{B_{(1)}^{2}}{2}+(\mn@boldsymbol{B_{(0)}}\!\cdot\!\mn@boldsymbol{B_{(1)}})+D\frac{(1-v^{2})^{1/2}}{\kappa}\,.$
Computations of the term (${S}\!\cdot\!\mn@boldsymbol{B})$ may also involve subtraction of large numbers and hence results in large errors. This can be avoided if we note that (${S_{(0)}}\!\cdot\!\mn@boldsymbol{B_{(0)}})=0$ and write
(${S}\!\cdot\!\mn@boldsymbol{B})=(\mn@boldsymbol{S_{(0)}}\!\cdot\!\mn@boldsymbol{B_{(1)}})+(\mn@boldsymbol{S_{(1)}}\!\cdot\!\mn@boldsymbol{B})\,.$
@@
$W^{2}v^{2}+4\bar{{\cal E}}_{1}W+4(p-W)\left(W+\frac{B^{2}}{2}\right)=S_{(1)}^{2}+2(\mn@boldsymbol{S_{(1)}}\!\cdot\!\mn@boldsymbol{S_{(0)}})-2\tilde{{\cal E}}_{1}B^{2}-B_{(0)}^{2}v_{0}^{2}(B_{(1)}^{2}+2(\mn@boldsymbol{B_{(0)}}\!\cdot\!\mn@boldsymbol{B_{(1)}}))\,,$ (152)
@@
~$\tilde{{\cal E}}_{1}={\cal E}_{(1)}-\frac{B_{(1)}^{2}}{2}-(\mn@boldsymbol{B_{(0)}}\!\cdot\!\mn@boldsymbol{B_{(1)}})\,.$
Appendix B Parameters of 1D shock simulations
B.1 Fast shock FS5
Left state:
${v}$ =( 0.99968283E+00, 0.00000000E+00, 0.00000000E+00)
${B}$ =( 0.50000000E+02, 0.19853866E+04, 0.00000000E+00)
${E}$ =( 0.00000000E+00, 0.00000000E+00, -0.19847569E+04)
p = 0.10000000E+01, $\rho$ = 0.10000000E+01, $\sigma$=1.E3.
Right state:
${v}$ = ( 0.99768146E+00, 0.17248747E-01, -0.00000000E+00)
${B}$ =( 0.50000000E+02, 0.19886156E+04, 0.00000000E+00)
${E}$ =( -0.00000000E+00 , 0.00000000E+00, -0.19831425E+04)
p = 0.44243911E+01, $\rho$ = 0.26176303E+01.
Shock speed $v_{s}=-0.5$, Mach number $M_{f}$=2.0.
The domain is $(-5.5,0.5)$ with 300 grid points. Initially, the shock is located at $x=0$.
B.2 Fast shock FS5A (shock FS5 in the rest frame of its left state)
Left state:
${v}$ = ( 0.00000000E+00 0.00000000E+00 0.00000000E+00)
${B}$ = ( 0.50000000E+02 0.50000000E+02 0.00000000E+00)
${E}$ = ( 0.00000000E+00 0.00000000E+00 0.00000000E+00)
p = 0.10000000E+01, $\rho$ = 0.10000000E+01
Right state:
${v}$ = ( -0.75954175E+00 0.16485693E+00 -0.00000000E+00)
${B}$ = ( 0.50000000E+02 0.24230060E+03 0.00000000E+00)
${E}$ = ( -0.00000000E+00 0.00000000E+00 0.19228027E+03)
p = 0.44243911E+01, $\rho$ = 0.26176303E+01
Shock speed $v_{s}$=-0.99989427E+00, Mach number $M_{f}$=2.0.
The domain is $(-1.5,0.5)$ with 100 grid points. Initially, the shock is located at $x=0$.
B.3 Fast shock FS7
Left state:
${v}$ =( 0.57368310E+00, 0.00000000E+00, 0.00000000E+00)
${B}$ =( 0.22803509E-01, 0.27840482E-01, 0.00000000E+00)
${E}$ =( 0.00000000E+00, 0.00000000E+00 , -0.15971614E-01)
p = 0.10000000E-01, $\rho$ = 0.10000000E+01, $\sigma$=0.001
Right state:
${v}$ = ( 0.19727530E+00, 0.34774998E-02, -0.00000000E+00)
${B}$ =( 0.22803509E-01, 0.13638473E+00, 0.00000000E+00)
${E}$ =(0.00000000E+00, 0.00000000E+00, -0.26826038E-01)
p = 0.28341867E+00, $\rho$ = 0.58282475E+01.
Shock speed $v_{s}=0.1$, Mach number $M_{f}$=5.0.
The domain is $(-0.5,1.5)$ with 100 grid points. Initially, the shock is located at $x=0$.
B.4 Fast shock FS9
Left state:
${v}$ =( 0.00000000E+00, 0.00000000E+00, 0.00000000E+00)
${B}$ =( 0.70724819E-02, 0.70724819E-02, 0.00000000E+00)
${E}$ =( 0.00000000E+00, 0.00000000E+00, 0.00000000E+00)
p = 0.10000000E-03, $\rho$ = 0.10000000E+01, $\sigma$=0.0001
Right state:
${v}$ = ( -0.57243160E-01, 0.31963724E-02, -0.00000000E+00)
${B}$ =(0.70724819E-02, 0.39243124E-01, 0.00000000E+00)
${E}$ =(0.00000000E+00, 0.00000000E+00, 0.22690067E-02)
p = 0.34131551E-02, $\rho$ = 0.52994146E+01.
Shock speed $v_{s}=-0.70530352E-01$, Mach number $M_{f}$=5.0.
The domain is $(-0.35,0.05)$ with 100 grid points. Initially, the shock is located at $x=0$.
B.5 Slow shock SS1
Left state:
${v}$ =( 0.19953950E+00 0.00000000E+00 0.00000000E+00)
${B}$ =( 0.50000000E+02 0.51026147E+02 0.00000000E+00)
${E}$ =( 0.00000000E+00 $0$0.10181732E+02)
p = 0.10000000E+01, $\rho$ = 0.10000000E+01, $\sigma$=1.E3.
Right state:
${v}$ = ( -0.42122856E+00 -0.63382468E+00 -0.00000000E+00)
${B}$ =( 0.50000000E+02 0.50825161E+02 -0.00000000E+00)
${E}$ =( 0.00000000E+00 $0$0.10282225E+02)
p = 0.14412306E+02, $\rho$ = 0.58792375E+01.
Shock speed $v_{s}=-0.5$, Mach number $M_{s}$=2.10.
The domain is $(-1.5,0.5)$ with 100 grid points. Initially, the shock is located at $x=0$.

### v3

@@ Appendix B Variables conversion
${\cal E}_{(1)}={\cal E}-{\cal E}_{(0)}=\mn@boldsymbol{E}_{(0)}\cdot\mn@boldsymbol{E}_{(1)}+\mn@boldsymbol{B}_{(0)}\!\cdot\!\mn@boldsymbol{B}_{(1)}+\frac{E_{(1)}^{2}+B_{(1)}^{2}}{2}+w\gamma^{2}-p\,,$ (138)
@@
${S}_{(1)}=\mn@boldsymbol{S}-\mn@boldsymbol{S}_{(0)}=\mn@boldsymbol{E}_{(0)}\!\times\!\mn@boldsymbol{B}_{(1)}+\mn@boldsymbol{E}_{(1)}\!\times\!\mn@boldsymbol{B}_{(0)}+\mn@boldsymbol{E}_{(1)}\!\times\!\mn@boldsymbol{B}_{(1)}+w\gamma^{2}\mn@boldsymbol{v}\,,$ (141)
@@
${S}_{(0)}=\mn@boldsymbol{E}_{(0)}\!\times\!\mn@boldsymbol{B}_{(0)}\,.$ (143)
@@
${E}_{(1)}=-\mn@boldsymbol{E}_{(0)}-\mn@boldsymbol{v}\!\times\!\mn@boldsymbol{B}\,.$ (145)
@@
Next, we use the perfect conductivity condition (145) to eliminate ${E}_{(1)}$ from the expression (138) for ${\cal E}_{(1)}$. To this end, we first find that
${E}_{(0)}\!\cdot\!\mn@boldsymbol{E}_{(1)}=-E_{(0)}^{2}-\mn@boldsymbol{E}_{(0)}\!\cdot\!(\mn@boldsymbol{v}\!\times\!\mn@boldsymbol{B})\,,$
@@
$E_{(1)}^{2}=E_{(0)}^{2}+2\mn@boldsymbol{E}_{(0)}\!\cdot\!(\mn@boldsymbol{v}\!\times\!\mn@boldsymbol{B})+||\mn@boldsymbol{v}\!\times\!\mn@boldsymbol{B}||^{2}\,,$
@@
${E}_{(0)}\!\cdot\!\mn@boldsymbol{E}_{(1)}+\frac{E_{(1)}^{2}}{2}=-\frac{E_{(0)}^{2}}{2}+||\mn@boldsymbol{v}\!\times\!\mn@boldsymbol{B}||^{2}\,.$
@@
${\cal E}_{(1)}=\frac{1}{2}(B^{2}v^{2}-(\mn@boldsymbol{v}\!\cdot\!\mn@boldsymbol{B})^{2})+W-p-\frac{E_{(0)}^{2}}{2}+\frac{B_{(1)}^{2}}{2}+\mn@boldsymbol{B}_{(0)}\!\cdot\!\mn@boldsymbol{B}_{(1)}\,.$ (151)
@@
¯$\bar{{\cal E}}_{1}={\cal E}_{(1)}+\frac{E_{(0)}^{2}}{2}-\frac{B_{(1)}^{2}}{2}-\mn@boldsymbol{B}_{(0)}\!\cdot\!\mn@boldsymbol{B}_{(1)}\,,$ (152)
@@
$a_{2}=\frac{1}{2}(B^{2}v^{2}-E_{(0)}^{2})-{\cal E}_{(1)}+\frac{B_{(1)}^{2}}{2}+\mn@boldsymbol{B}_{(0)}\!\cdot\!\mn@boldsymbol{B}_{(1)}+D\frac{(1-v^{2})^{1/2}}{\kappa}\,.$
The first two terms of this expression constitute the difference between $B^{2}v^{2}/2$ and $E_{(0)}^{2}$. These non-negative terms can be very large and their difference can be a source of large error in computations of $a_{2}$ in the case of high magnetisation.
@@
${v}_{(0)}=\frac{\mn@boldsymbol{E}_{(0)}\!\times\!\mn@boldsymbol{B}_{(0)}}{B_{(0)}^{2}}\,.$
@@
$B^{2}v^{2}-E_{(0)}^{2}=B^{2}v^{2}-B_{(0)}^{2}v_{(0)}^{2}=B_{(0)}^{2}(v^{2}-v_{(0)}^{2})+(B_{(1)}^{2}+\mn@boldsymbol{B}_{(0)}\!\cdot\!\mn@boldsymbol{B}_{(1)})v^{2}\,,$
@@
$a_{2}$ $=\frac{1}{2}(B_{(0)}^{2}(v^{2}-v_{(0)}^{2})+(B_{(1)}^{2}+\mn@boldsymbol{B}_{(0)}\!\cdot\!\mn@boldsymbol{B}_{(1)})v^{2})-$
$-{\cal E}_{(1)}+\frac{B_{(1)}^{2}}{2}+\mn@boldsymbol{B}_{(0)}\!\cdot\!\mn@boldsymbol{B}_{(1)}+D\frac{(1-v^{2})^{1/2}}{\kappa}\,.$
Computations of the term ${S}\!\cdot\!\mn@boldsymbol{B}$ may also involve subtraction of large numbers and hence results in large errors. This can be avoided if we note that ${S}_{(0)}\!\cdot\!\mn@boldsymbol{B}_{(0)}=0$ and write
${S}\!\cdot\!\mn@boldsymbol{B}=\mn@boldsymbol{S}_{(0)}\!\cdot\!\mn@boldsymbol{B}_{(1)}+\mn@boldsymbol{S}_{(1)}\!\cdot\!\mn@boldsymbol{B}\,.$
@@
$W^{2}v^{2}$ $+4\bar{{\cal E}}_{1}W+4(p-W)\left(W+\frac{B^{2}}{2}\right)=S_{(1)}^{2}+2\mn@boldsymbol{S}_{(1)}\!\cdot\!\mn@boldsymbol{S}_{(0)}-$
$-2\tilde{{\cal E}}_{1}B^{2}-B_{(0)}^{2}v_{(0)}^{2}(B_{(1)}^{2}+2\mn@boldsymbol{B}_{(0)}\!\cdot\!\mn@boldsymbol{B}_{(1)})\,,$ (159)
@@
~$\tilde{{\cal E}}_{1}={\cal E}_{(1)}-\frac{B_{(1)}^{2}}{2}-\mn@boldsymbol{B}_{(0)}\!\cdot\!\mn@boldsymbol{B}_{(1)}\,.$